OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. x-intercept 5; y-intercept −10

myininaya (myininaya):

y=mx+b is the equation for a line x intercept is (5,0) y intercept is (0,-10) so we know two points on the line we can find m (the slope) slope=m=\[\frac{ \Delta y}{\Delta x}=\frac{0-(-10)}{5-0}=\frac{10}{5}=2\] so we have y=2x+b now using either point we can find b (the y-intercept) we have (x=5,y=0) so we have 0=2*5+b 0=10+b -10=b so the equation of the line that have intercepts (5,0) and (0,-10) is y=2x+(-10) or you can say y=2x-10

OpenStudy (anonymous):

what about this: Find an equation of the line that satisfies the given conditions. Through (2, 9); parallel to the line passing through (3, 7) and (−1, 3)

myininaya (myininaya):

find slope of the line containing points (3,7) and (-1,3) parallel lines have the same slope so whatever you find that slope to be use it for y=mx+b (then you can b by pluggin' in (2,9) ) let me know what you get i will check k

OpenStudy (anonymous):

im doin it wrong do you mind walkin me through it

myininaya (myininaya):

\[m=\frac{7-3}{3-(-1)}=\frac{4}{3+1}=\frac{4}{4}=1\] did you get this for slope?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so we have y=1x+b we know a point on the line (2,9) 9=1*2+b solve for b 9=2+b 9-2=b 7=b so the equation of the line going through (2,9) and is parallel to the line containing points (3,7) and (-1,3) is y=1x+7 or you can simply write y=x+7

OpenStudy (anonymous):

thanks so much